four identical cells each of EMF 1.5v and internal resistance of 0.5 ohms are connected in series across an external load of resistance 1.5 ohms. calculate the current that passes through the load


1
Expert's answer
2021-12-09T10:07:29-0500

V=IR

"I=\frac{V}{R}"

"R_T=R_i+R_L=0.5+1.5=2"

"I=\frac{1.5}{2}=0.75A"